OpenStudy (anonymous):

If a customer buys 3 wallet-sized prints for $1.99 each (total price including tax), how much change will she get from a $20 bill. $13.43 $15.00 $14.03 Not Here $5.97

OpenStudy (hoblos):

20 - 3(1.99) = 20 - 5.97 = $14.03

OpenStudy (anonymous):

thank you :D